OpenStudy (anonymous):

The volume of a cylinder is 289 cm2 and the height is 21 cm. Find the diameter.

OpenStudy (aaronq):

\[V _{cylinder}=\pi*r ^{2}*h\] \[r=\frac{ 1 }{ 2 }d\] \[V _{cylinder}=\pi* (\frac{ 1 }{ 2 }d)^{2}*h\] \[d=2*\sqrt{\frac{ V _{cylinder} }{ h*\pi }}\]

OpenStudy (aaronq):

plug in the height (h) and volume (V) given.